For the graph shown, select the statement that best represents the given system of equations.
algebra unit 8 lesson 1 assessment question 4
2y – 3x = 9
4y + 3x = 12

A.
consistent and independent

B.
inconsistent

C.
coincident

D.
not enough information

Answer by TimothyLamb(4379)

2y – 3x = 9
4y + 3x = 12
---
put the system of linear equations into standard form
---
-3x + 2y = 9
3x + 4y = 12
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -0.66666667
y = 3.5
---
answer:
the linear system is: A. consistent and independent
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php